OpenStudy (anonymous):

A deck of 52 cards is mixed well, and 5 cards are dealt.What is the probability that a hand will consist entirely of a single suit?

OpenStudy (welshfella):

taking 1 card at a time P(say Spade) = 1/4 ( because there are 13 spades out of 52) 2nd card P(spade) = 12/51 3rd card = 11/50 so can you tell me the next 2 probabilities of getting a spade?

OpenStudy (anonymous):

10/49 and 9/48??

OpenStudy (welshfella):

right now as all these probabilities are independent to get P(5 spades) you multiply them all together

OpenStudy (welshfella):

1/4 * 12/51 * 11/50 * 10/49 * 9/48